What does 4[(6-1)+3(5-2)] Equal?
I keep getting 49 why?

Found 3 solutions by stanbon, jim_thompson5910, Alan3354:
Answer by stanbon(75887)

4[(6-1)+3(5-2)]
-------------------
4[5+3*3]
---
= 4[5+9]
= 4*14
= 56
==============
Cheers,
Stan H.

Answer by jim_thompson5910(35256)

4[(6-1)+3(5-2)]

4[(5)+3(3)]

4[(5)+9]

4[5+9]

4[14]

56

So 4[(6-1)+3(5-2)] = 56

Answer by Alan3354(69443)

What does 4[(6-1)+3(5-2)] Equal?
4[(6-1)+3(5-2)] = 4[5 + 3*3]
= 4[5 + 9]
= 4*14
= 56